OpenStudy (anonymous):

Use the Pythagorean theorem to find the length of the hypotenuse with one side measuring 16 feet and the other measuring 7 feet.

OpenStudy (anonymous):

The Pythagorean theorem states : a² + b² = h² with a,b length of the adjacent sides to the right angle, and h length of the hypotenuse. Then, you got : 16² + 7² = h² which you can simply solve.

OpenStudy (anonymous):

\[c ^{2} = a^2 + b^2 \] \[c ^{2} = 16^2 + 7^2 \] \[c ^{2} = 256 + 49 \] \[c ^{2} = 301\] \[c =\sqrt[2]{30}\]

OpenStudy (anonymous):

\[c = \sqrt{301}\] i mean.

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] \[16^2 + 7^2 = c\] \[256 + 49 = c^2\] \[305 = c^2\] \[\sqrt{305} = c\] \[17.5 \approx c\]

OpenStudy (anonymous):

@josephniel: You added 256 + 49 wrong o_0

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Oh right. Sorry.